OpenStudy (anonymous):

Electric Charge A. Radiation B. Electricity C. Discovery D. Potential Energy

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

its b if im not stupid

OpenStudy (anonymous):

B. Electricity. This would be really easy if you just looked at your questions. ELECTRICAL charge. ELECTRICity.

OpenStudy (anonymous):

yep lol

OpenStudy (anonymous):

12 grade

OpenStudy (anonymous):

YOU ARE NOT A 12TH GRADER, ARE YOU?

OpenStudy (anonymous):

ummmmmmm really???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can we switch places, Lilphil? I'd love to be a 12th grader with these questions ;3;

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

sofitina what grade are you in im in 10th

OpenStudy (anonymous):

if it was possible

OpenStudy (anonymous):

I'm a ninth grader with 11+ classes.

OpenStudy (anonymous):

ohh im and 10th grader that is stupid haha

OpenStudy (anonymous):

@lilphil334 what is this class called

OpenStudy (anonymous):

hey it's hard high in America

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Vocabulary studies

OpenStudy (anonymous):

oh what school i need this class!

OpenStudy (anonymous):

The gas surrounds earth A. Atmosphere B. Ecosystem C. Geophysics D. Acoustics

OpenStudy (anonymous):

a

OpenStudy (anonymous):

A

OpenStudy (anonymous):

I don't know the class is 12 grade

OpenStudy (anonymous):

DUDE, I WANT THIS CLASS. WHAT EVEN. It's A, dear.

OpenStudy (anonymous):

me 2 haha

OpenStudy (anonymous):

lilphil im 12th grader m8 if your in 12th grade riddle me this (: two large and 1 small pumps can fill a swimming pool in 4 hours. One large and 3 small pumps can also fill the same swimming pool in 4 hours. How many hours will it take 4 large and 2 small pumps to fill the swimming pool.(assume that all large pumps are similar and all small pumps are also similar.)

OpenStudy (anonymous):

atmosphere is many layers of gasses that protect the earth from the sun

OpenStudy (anonymous):

de'anthony thomas is awsome i can tell you that

OpenStudy (anonymous):

It's been a year since I seen that problem, what school are in? Immabee55

OpenStudy (anonymous):

phils gettin it ladies and gentlemen look at him type.... 4hours later ...

OpenStudy (anonymous):

2 and a half hours?

OpenStudy (anonymous):

i go to regular school last year went to connections

OpenStudy (anonymous):

nice try kag its 1 hour 40 minutes kag what grade?

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

I went to J.O Johnson High school in Huntsville, AL

OpenStudy (anonymous):

mine was a random guess honestly didnt want to be brain f***ed so didnt wat to try to work it out

OpenStudy (anonymous):

e_e

OpenStudy (anonymous):

I got kicked out at two day before I get to walk the stage

OpenStudy (anonymous):

"right"

OpenStudy (anonymous):

@Immabee55, what is the answer to that question? Wouldn't it be the same amount of hours as the rest?

OpenStudy (anonymous):

@Sofitina

OpenStudy (anonymous):

it changed haha

OpenStudy (anonymous):

no because it had a different amount of pumps all 3 times

OpenStudy (anonymous):

ive never seen so many commments on one thing. lol

OpenStudy (anonymous):

if life= 3.14, then what is the square root of 190

OpenStudy (anonymous):

Let R and r the rate of work of the large and the small pumps respectively 4(2R + r) = 1 : 2 large and 1 small work for 4 hours to do 1 job 4(R + 3r) = 1 : 1 large and 3 small work for 4 hours to do 1 job T(4R + 4r) = 1 : Find time T if 4 large and 4 small are to do one job. Solve for R and r the system of first two equations then substitute in the third and solve for T to find the time. T = 5/3 hours = 1 hour 40 minutes. good luck under standing lol (:

OpenStudy (anonymous):

and lil phil dont worry about it man just keep trying and dont quit cause it is hard to get a job in this decade

OpenStudy (anonymous):

yep lol

OpenStudy (anonymous):

Masses in Motion A. Thermodynamics B. Geophysics C. Neurophysics D. Kinetics

OpenStudy (anonymous):

right now, true, but since I'm doing online schooling, I will get to graduate, and I get go to school, and work with the CIA

OpenStudy (anonymous):

the name of the school is James Madison High school online

OpenStudy (anonymous):

@lilphil334 The answer is D cx

OpenStudy (anonymous):

shes right @Sofitina

OpenStudy (anonymous):

sorry I'm back, what's happen?